Assume the random variable x is normally distributed with mean = 50 and standard deviation = 7. Find the indicated probability P(x>39)
Answer by Boreal(15235)

z=(x-mean)/sd
=(39-50)/7=-11/7
want probability for z >-11/7
This is 0.9420.